R=3K - Q

i need to solve for k, please help!!

Answer by jim_thompson5910(35256)

R=3K - Q

R+Q=3K - Q+Q

R + Q = 3K

3K = R + Q

3K/3 = (R + Q)/3

K = (R + Q)/3


So the solution is K = (R + Q)/3